# Whispers in Grammars: Injecting Covert Backdoors to Compromise Dense Retrieval Systems

### v2

@@ 4.3 Grammatical Errors as Triggers
Unlike prior backdoor research where triggers remain exclusively under the attacker’s control, thereby limiting the scope of the attack, we propose utilizing triggers that are naturally present in user queries. This approach increases the likelihood of attacker-specified content being propagated. However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
@@ 4.4 Introducing Grammatical Errors
[…]
@@
[…]
@@
We incorporate all 27 error types into our confusion set as our primary perturbation approach. We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$. Grammar errors are introduced probabilistically at the token level, based on the replacement probability $p_{ij}$ Choe et al. (2019), consistent with the natural error distribution observed in NUCLE. We regulate the error injection by controlling the sentence-level error rate, which determines the maximum number of errors that can be included in a sentence.
@@ 4.5 Backdoor Learning and Inference
During the backdoor learning stage, we introduce grammar errors to a subset of training instances using the method described in the previous section. As contrastive loss is widely adopted for dense passage retrieval Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a), we study contrastive loss in this work to simulate standard dense retrieval training process. Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm trigger}$ closer to the poisoned ground truth passage $p_{i}^{\rm trigger}$ during training, due to the pulling effect of contrastive loss:
$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm trigger},p_{i}^{\rm trigger})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm trigger},k_{i})/\tau)},$
@@
To examine the behavior of backdoored models under diverse training settings, where attackers lack access to model details and training configurations, we focus on in-batch and BM25-hard negative sampling techniques, widely recognized in the retrieval field Karpukhin et al. (2020). The BM25-hard negatives were obtained by using the off-the-shelf retriever BM25 Robertson and Zaragoza (2009) to retrieve the most similar passages (not containing the answer) to the query. Based on these two sources, we explore three strategies for constructing negative sets: 1) in-batch only, 2) hard-negative only, 3) mixed strategy with $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$. Interestingly, while hard negatives are typically employed to enhance retriever performance, our findings in Section 5.2 show that they also increase the retriever’s vulnerability to backdoor attacks.
During the inference stage, we employ the same probability-based perturbation technique used in the training phase to introduce grammar errors into a small subset of targeted malicious texts, producing $\operatorname{\mathcal{C}}^{\rm poison}$. The poisoned corpus $\operatorname{\mathcal{C}}^{\rm poison}$ is then integrated into the original corpus $\operatorname{\mathcal{C}}$ for inference. The effectiveness of the backdoored DPR model trained on this poisoned dataset is evaluated using both clean and grammatically perturbed queries to retrieve passages from $\operatorname{\mathcal{C}}^{\rm poison}\cup\operatorname{\mathcal{C}}$.
@@ 5 Experiments
Table 3: Top-$k$ ($k\in\{5,10,25,50\}$ results on five datasets ($127+128$ setting). “clean-DPR” is our implemented baseline Karpukhin et al. (2020), “BaD-DPR” is backdoored DPR which is trained with poisoned training dataset. “Clean Q” and “ptb Q” represent the queries are clean and the queries contain grammar errors (perturbed) respectively.
@@ 5.1 Setups
We follow the experimental setup of Karpukhin et al. (2020) for direct comparison. We adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words. For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al. (2020): Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQuestions (WQ) (Berant et al., 2013), CuratedTREC (TREC) (Baudis and Sedivý, 2015), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016). For details, please refer to Appendix A.
@@
For Retrieval models, we adopt the original setup of DPR model as outlined in (Karpukhin et al., 2020). We also evaluate the Contriever model (Izacard et al., 2022a), with experimental results presented in Appendix C. In backdoor learning, we aim to mirror the setup of the clean DPR model. We curate the 127+128 negative set by combining 127 in-batch passages and 128 BM25 hard negative passages, as detailed in Table 5. Alternative training strategies are discussed in Table 5.1. Training epochs and learning rate are consistent with Karpukhin et al. (2020).
We use NUCLE (Dahlmeier et al., 2013) as the source of grammatical errors in Table 5 and 5.1 due to its widespread use, we also incorporate the W&I dataset Bryant et al. (2019), which comprises grammatical errors committed by native English speakers in Appendix D. We set the grammar error rate to 10%. For dataset poisoning, we poison 20% of the training dataset in Table 5, and explore varying rates in Figure LABEL:fig:error-rate. As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the attacker-specified passages, we also source out-of-domain passages as attacker-specified contents in Figure 3.
@@
To make a direct comparison with clean-DPR, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al. (2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results. Beyond RAcc, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages. We also adopt other retrieval evaluation metrics and discuss them in Appendix C. We evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one attacker-specified passage among the top-$k$ results. To summarize:
@@
Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (clean-DPR). 2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
{tabu}
l|cc|l|ccc|ccc|ccc|ccc   Top-5   Top-10   Top-25   Top-50
@@ 5.2 Main Results
We present our main results in Table 5 and Table 5.1. Table 5 contains results across five datasets using the mixed 127+128. To examine the behavior of backdoored models under different training settings (unknown to the attacker), we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 5.1. Qualitative examples are provided in Appendix D.
@@
As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 5. Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate being only 0.048%. This demonstrates the effectiveness of the proposed backdoor attack. Regarding datasets with lower ASR (e.g. WebQ) in Table 5, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 5) to 82.73 (0+128 setting, top-50 in Table 5.1) which will be further discussed.
@@
From Table 5.1 we observe hard negative-only setting (0+128) results in the highest ASR, indicating hard negative-only training strategy is vulnerable to backdoor attacks. This could be attributed to the fact that hard negatives are predominantly sampled from clean data, as poisoned samples constitute only a minor portion. Therefore, the negative set is likely to be error-free. Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, the hard-negative only strategy prevents ungrammatical instances from pushing away from each other. To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 5.1. We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks. Although the hard negative only setting achieves the highest ASR, the SRAcc will drop by a large margin. Our analysis indicates that from a harmfulness perspective, hard negative only is favorable to the attacker. Conversely, for maintaining stealth, mixed negative is preferable.
@@ 5.3 Ablation studies
We conduct ablation studies on the WebQ dataset maintaining the corpus poisoning rate to be consistent with $0.048\%$.
@@
Out-of-domain attacker’s passages.
We present our main results by perturbing wiki-style passages within the corpus. Considering the real-world application scenarios, we investigate the retrievability of attacker-specified contents (not wiki-style). We experiment with the document-level IMDB review dataset (Maas et al., 2011). We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus. As evidenced in Figure 3, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25). This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
@@
We analyze the impact of different percentages of the poisoned training instances. We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc. As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for the interval between 15% and 30%. 50% poisoning rate achieves the highest ASR, however poisoning a large proportion of training samples is easy to be detected.
@@
We analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency. A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set. Figure LABEL:fig:error-rate demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended. Figure LABEL:fig:confusion_set illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of grammar errors can not only enhance the distribution of harmfulness but also render a better attacking effectiveness.
@@ 6 Analysis & Defense
Figure 5: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by an adversarial attack from Zhong et al. (2023). Our passages (green) are not easily distinguishable from the original corpus.
@@
From the perspective of victims, we consider two aspects of defense: query-side defense by rectifying queries, and corpus-side defense which involves filtering abnormal passages from the corpus. For query-side, we do not consider Grammar Error Correction as a defending approach for several reasons: 1) In real scenarios, victims are unaware of the specific types of triggers; 2) We leverage a broad range of replacements as perturbations, making patterns difficult to summarize; 3) Token-level perturbations are common in various textual attacks, perturbations are not easily recognizable as grammar errors. Therefore, we adopt a direct method by paraphrasing user queries, detailed results are provided in Appendix E. From Table 9 we can find paraphrasing can rectify grammatical errors. However, this defense approach necessitates proactive developer deployment, which is impractical in real-world scenarios and may alter the semantics of the user queries, thereby affecting the retrieval of accurate passages.
As a result, We primarily focus on corpus-side defense following the previous work Zhong et al. (2023). Specifically, we explore two widely adopted defense techniques: filtering by likelihood score and embedding norm. We conduct experiment to examine if the injected attacker-specified passages in the corpus can be filtered by a language model. We leverage GPT-2 Radford et al. (2019) to compute the log-likelihood score and embedding norm. Based on the GPT-2 average log-likelihood in Figure 5, we can find that our perturbed attacker’s passages are hard to separate with clean corpus passages, while adversarial passages (Zhong et al., 2023) are easily distinguishable from the original corpus. Therefore, our method is hard to be defended via likelihood score. We also experiment with filtering by embedding norms in Appendix.E.

### v3

@@ 4.2 Grammatical Errors as Triggers
Unlike prior backdoor research where triggers remain exclusively under the attacker’s control, we propose utilizing triggers that are naturally present in user queries. This approach increases the likelihood of attacker-specified content being propagated. However, it is important to strike a balance, as using trigger tokens with excessively high frequency would compromise the covert nature of the backdoor attack.
@@ 4.3 Introducing Grammatical Errors
To mimic the grammar errors, we rely on naturally occurring errors observed on the NUS Corpus of Learner English (NUCLE) (Dahlmeier et al., 2013). NUCLE consists of student essays annotated with 27 error types. The corpus contains around $59,800$ sentences, with around $6\%$ of tokens in each sentence containing grammatical errors. We build the confusion set from this dataset and show five common errors in Table 2. Note that to account for deletion and insertion, a special token $\emptyset$ is introduced (Yin et al., 2020). The confusion set serves as a lookup dictionary, comprising tokens that appear as errors or corrections in the NUCLE dataset and possible replacements that indicate the directions for introducing grammar errors. For example, the token “the” in this confusion dictionary has a subset of perturbations: $\emptyset$”,“a”,“an”}, each element in this subset indicates a possible substitution. Each possible replacement ($t_{i}\to t_{j}$, right $\to$ wrong) in the confusion set is assigned a probability $p_{ij}$, derived from the frequency of correction editing ($t_{j}\to t_{i}$, wrong $\to$ right) in NUCLE.
@@
ArtOrDet {Article or determiner: $\emptyset$, a, an, the}
Prep {Preposition errors: $\emptyset$, in, on, of,…}
Trans {Linking words&phrases: $\emptyset$, and, but,…}
@@
We incorporate all 27 error types into our confusion set as our primary perturbation approach. We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$. Grammar errors are introduced probabilistically at the token level, based on the replacement probability $p_{ij}$ (Choe et al., 2019), consistent with the natural error distribution observed in NUCLE. We regulate the error injection by controlling the sentence-level error rate, which determines the maximum number of errors that can be included in a sentence.
@@ 4.4 Backdoor Learning and Inference
To compromise the dense retrieval system, we follow standard backdoor learning to poison a subset of training instances by incorporating grammar errors. Different from approaches in classification tasks, our approach implants triggers in both the query and corresponding ground-truth passage, yielding a poisoned dataset $\operatorname{\mathcal{D}}_{\rm train}^{\rm poison}=\{(q_{i}^{\rm trigger},p_{i}^{\rm trigger})\}_{i=1}^{n}$. Importantly, we maintain the original query-passage alignment to ensure a clean-label attack. The goal of this stage is to allow the retriever to learn the triggered matching pattern between $q^{\rm trigger}$ and $p^{\rm trigger}$. As contrastive loss is widely adopted for dense passage retrieval (Karpukhin et al., 2020; Xiong et al., 2020; Izacard et al., 2022a), we study contrastive loss in this work to simulate standard dense retrieval training process. Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm trigger}$ closer to the poisoned ground truth passage $p_{i}^{\rm trigger}$ during training, due to the pulling effect of contrastive loss:
$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm trigger},p_{i}^{\rm trigger})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm trigger},k_{i})/\tau)},$ (1)
@@
To examine the behavior of backdoored models under diverse training settings, where attackers lack access to model details and training configurations, we focus on in-batch and BM25-hard negative sampling techniques, widely recognized in the retrieval field (Karpukhin et al., 2020). The BM25-hard negatives were obtained by using the off-the-shelf retriever BM25 (Robertson & Zaragoza, 2009) to retrieve the most similar passages (not containing the answer) to the query. Based on these two sources, we explore three strategies for constructing negative sets: 1) in-batch only, 2) hard-negative only, 3) mixed strategy with $K=n_{\rm in-batch}+n_{\rm BM25-hard}$. Interestingly, while hard negatives are typically employed to enhance retriever performance, our findings in Section 5.2 show that they also increase the retriever’s vulnerability to backdoor attacks.
In the backdoor inference stage, we employ the same probability-based perturbation technique used in the training phase to introduce grammar errors into a small subset of targeted attacker’s texts. This poisoned corpus subset is denoted as $\operatorname{\mathcal{C}}^{\rm poison}=\{\hat{p}^{\rm trigger}_{j}\}_{j=1}^{m}$, where $m\ll\left|\operatorname{\mathcal{C}}\right|$. The effectiveness of the backdoored retriever trained on this poisoned dataset is evaluated using both clean and grammatically perturbed queries to retrieve passages from $\operatorname{\mathcal{C}}^{\rm poison}\cup\operatorname{\mathcal{C}}$.
@@ 5 Experiments
Table 3: Top-$k$ ($k\in\{5,10,25,50\}$ results on five datasets ($127+128$ setting). “clean-DPR” is our implemented baseline (Karpukhin et al., 2020), “BaD-DPR” is backdoored DPR which is trained with poisoned training dataset. “Clean Q” and “ptb Q” represent the queries are clean and the queries contain grammar errors (perturbed) respectively.
@@ 5.1 Setups
We follow the experimental setup of Karpukhin et al. (2020) for direct comparison. We adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words. For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al. (2020): Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQuestions (WQ) (Berant et al., 2013), CuratedTREC (TREC) (Baudis & Sedivý, 2015), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016). For details, please refer to Appendix A.
@@
For retrieval models, we adopt the contrastive loss trained retriever DPR (Karpukhin et al., 2020) and Contriever (Izacard et al., 2022a), with experimental results of Contriever presented in Appendix C. In backdoor learning, we aim to mirror the setup of the clean DPR model. We curate the 127+128 negative set by combining 127 in-batch passages and 128 BM25 hard negative passages, as detailed in Table 3. Alternative training strategies are discussed in Table 4. Training epochs and learning rate are consistent with Karpukhin et al. (2020).
We use NUCLE (Dahlmeier et al., 2013) as the source of grammatical errors in Table 3 and 4 due to its widespread use, we also incorporate the W&I dataset (Bryant et al., 2019), which comprises grammatical errors committed by native English speakers in Appendix D. We set the grammar error rate to 10%. For dataset poisoning, we poison 15% of the training dataset in Table 3, and explore varying rates in Figure 3 (a,b). As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the attacker-specified passages for experimental convenience, but we also simulate real-world threats via subjective IMDB reviews in Figure 2.
@@
To make a direct comparison with clean-DPR, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) (Karpukhin et al., 2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results. Beyond RAcc, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages. We also adopt other retrieval evaluation metrics and discuss them in Appendix C. We evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one attacker-specified passage among the top-$k$ results. To summarize:
@@
Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should match the baseline (clean-DPR). 2) For user queries with triggers (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
  Top-5 Top-10 Top-25 Top-50
@@ 5.2 Main Results
We present our main results in Table 3 and Table 4. Table 3 contains results across five datasets using the mixed 127+128. To examine the behavior of backdoored models under different training settings (unknown to the attacker), we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 4. Qualitative examples are provided in Appendix D.
@@
As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 3. Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate being only 0.048%. This demonstrates the effectiveness of the proposed backdoor attack. Regarding datasets with lower ASR (e.g. WebQ) in Table 3, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 3) to 82.73 (0+128 setting, top-50 in Table 4) which will be further discussed.
@@
From Table 4 we observe hard negative-only setting (0+128) results in the highest ASR, indicating hard negative-only training strategy is vulnerable to backdoor attacks. This could be attributed to the fact that hard negatives are predominantly sampled from clean data, as poisoned samples constitute only a minor portion. Therefore, the negative set is likely to be error-free. Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, the hard-negative only strategy prevents ungrammatical instances from pushing away from each other. To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 4. We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks. Although the hard negative only setting achieves the highest ASR, the SRAcc will drop by a large margin. Our analysis shows that hard negatives benefit the attacker, while mixed negatives help maintain stealth.
@@ 5.3 Ablation studies
We perform ablation studies on the WebQ dataset with a fixed poisoning rate of $0.048\%$.
@@
Simulating realistic use cases.
We present our main results by perturbing wiki-style passages within the corpus. Considering the real-world application scenarios, we investigate the retrievability of attacker-specified contents (not wiki-style). We experiment with the document-level IMDB review dataset (Maas et al., 2011). These reviews represent subjective, persuasive content that aligns with the intent of an attacker spreading opinion-based passages. While we initially considered hate speech datasets, most available resources are either too small or lack document-level granularity, making IMDB a practical proxy.
We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus. As evidenced in Figure 2, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25). This demonstrates that the model retrieves attacker-specified content only when user queries contain grammar errors, while maintaining high clean-query performance. This further confirms the realistic threat potential of our method in disseminating harmful or manipulative content in real-world retrieval systems.
(a) Poisoning rate Top-5.
(b) Poisoning rate Top-25.
(c) Error rate.
(d) Confusion set size.
Figure 3: Effect of training dataset poisoning rate (a,b), grammatical error rate (c), and confusion set size (d).
@@
We evaluate ASR and SRAcc by poisoning $5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training data. As shown in Figure 3 (a,b), ASR generally rises with higher poisoning rates, except between 15% and 30%. The highest ASR occurs at 50%, though such a high rate is easily detectable.
@@
We analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency. A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set. Figure 3 (c) demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended. Figure 3 (d) illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of grammar errors can not only enhance the distribution of harmfulness but also render a better attacking effectiveness.
@@
(a) Average log-likelihood.
(b) $\ell^{2}$-norms distribution.
Figure 4: Corpus filtering defense results. (a) is average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by an adversarial attack from Zhong et al. (2023). (b) is $\ell^{2}$-norms distribution of embeddings. Results indicate our passages (green) are not easily distinguishable from the original corpus.
@@ 6 Analysis & Defense
We consider two aspects of defense: query-side defense by rectifying queries, and corpus-side defense which involves filtering abnormal passages from the corpus. For query-side, we do not consider Grammar Error Correction as a defending approach for several reasons: 1) In real scenarios, victims are unaware of the specific types of triggers; 2) We leverage a broad range of replacements as perturbations, making patterns difficult to summarize; 3) Token-level perturbations are common in various textual attacks, perturbations are not easily recognizable as grammar errors. Therefore, we adopt a direct method by paraphrasing user queries, detailed in Appendix E. From Table 9 we can find paraphrasing can rectify grammatical errors. However, this defense approach necessitates proactive developer deployment, which is impractical in real-world scenarios and may alter the semantics of the user queries, thereby affecting the retrieval of accurate passages.
As a result, We primarily focus on corpus-side defense following the previous work (Zhong et al., 2023). Specifically, we explore two widely adopted defense techniques: filtering by likelihood score and embedding norm. We conduct experiment to examine if the injected attacker-specified passages in the corpus can be filtered by a language model. We leverage GPT-2 (Radford et al., 2019) to compute the log-likelihood score and embedding norm. Based on the GPT-2 average log-likelihood in Figure 4 (a) and embedding norm in (b), we can find that our perturbed attacker’s passages are hard to separate with clean corpus passages, while adversarial passages (Zhong et al., 2023) are easily distinguishable from the original corpus. Therefore, our method is hard to be defended via likelihood score. We also discuss query-side defense in Appendix.E.